night-report-page-20131102

Observers: Massimo Turatto, Susanna Spiro TO: Ariel Sanchez

Support team: S.Benetti, N.Elias-Rosa

TAT on duty: Maria Teresa Botticella, Annalisa De Cia

------------------------------------------------

Subrun 2, night 2

Instruments: SOFI

WEATHER CONDITION

beginning: wind speed 2 m/s; humidity 34 % ; despejado- photometric

middle: wind speed 3 m/s; humidity 30 %; clear

end: wind speed 2 m/s; humidity 10 %; clear

afternoon CALIBRATIONS

PESSTO_SOFI_SpecArc

PESSTO_SOFI_specflat

OBSERVATIONS: start regularly at sunset.

Target Type UT Filter/Grism Slit Seeing Grade (A/B/C) Comments

SN2012ca follow-up 23:30 JHKs 0.9 A

SN2013ek follow-up 00:55 J 0.8 A

SN2013ej follow-up 01:20 JHKs 1.3 A

LTT 7987 fluxSTD 01:35 BG,RG 1.0 A Single exposures should be longer

SN2013fc follow-up 01:50 JHKs 0.8 A

during Ks the probe gets stuck while offsetting. Second series of Ks aborted but SN well visible in the first.

Then the telescope got stuck. Reboot of the TCS.Lost ~1h.

Hip014339 tellSTD 04:27 BG, RG 1.0 0.7 A

SN2013fc follow-up 04:36 BG, RG 1.0 0.6 A visible also in single exposures

Hip016853 tellSTD 08:30 BG, RG 1.0 0.6 A

OGLE-2013-SN-079 follow-up 08:50 J 1.1 A still invisible